### v2

@@ 6 Conclusion
In this paper, I propose a new TERC model in which the regressors are correlated with the random coefficients not only through a fixed effect but, more importantly, through a time-varying random shock—a feature that aligns with agent optimization behavior in many empirical applications. I construct feasible control variables for both the fixed effect and the random shock, and subsequently utilize the residual variation in the regressors given the control variables to identify the APE and LAR. I provide three-step series estimators and establish their convergence rates and asymptotic normality. The empirical exercise of this paper reveals significant variation in the output elasticities across manufacturing firms in China.
I propose two directions for future research. First, beyond the first- and second-order moments analyzed in this paper, policymakers may also be interested in the density functions of the random coefficients. One possible, albeit laborious, approach is to identify the moments of all orders of the random coefficients by induction, which would uniquely determine their distribution (Stoyanov (2000)). Second, it remains an open question whether aspects of my method could be extended to dynamic linear or nonlinear panel data models, such as those discussed by Marx, Tamer, and Tang (2024) and Liu, Poirier, and Shiu (2024). Such an extension would likely require additional structure on how the lagged dependent variable correlates with the random coefficients.
Appendix A Proofs
@@
I first prove that $V_{it}$ is a control for $\eta_{it}$ given $A_{i}$ and $W_{i}$. Then, I show $\mathbb{E}\left[\left.\beta_{it}\right|X_{it},V_{it},W_{i}\right]$ does not depend on $X_{it}$ via the LIE. Finally, I identify $\mathbb{E}\beta_{it}$ and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ by leveraging the residual variation in $X_{it}$ given $V_{it}$ and $W_{i}$. I show how the inclusion of exogenous shocks $\upsilon_{it}$ and $\epsilon_{it}$ as well as exogenous coordinates in the regressors affects the analysis at the end of this proof.
Without loss of generality, I assume that $d_{\eta}=d_{X}$ and that each coordinate of $\eta_{it}$ enters the corresponding coordinate of $X_{it}$. [^26] This is without loss of generality because under Assumption 1, I can always redefine $\widetilde{\eta}$ to be the vector that collects the coordinates of $\eta$ that enters each coordinate of $g\left(\cdot\right)$ function. By Assumption 2, I have $\left.A_{i}\perp\left(X_{it},Z_{it}\right)\right|W_{i}$, which implies $\left.X_{it}\perp A_{i}\right|\left(Z_{it},W_{i}\right)$. Thus, for each $l\in\left\{1,...,d_{X}\right\}$ and any on-support $\left(x^{\left(l\right)},z,a,w\right)$, I have
$\ F_{\left.X_{it}^{\left(l\right)}\right|Z_{it},W_{i}}\left(\left.x^{\left(l\right)}\right|z,w\right)$
$=$ $\ F_{\left.X_{it}^{\left(l\right)}\right|Z_{it},A_{i},W_{i}}\left(\left.x^{\left(l\right)}\right|z,a,w\right)$
$=$ $\ \mathbb{P}\left(\left.g^{\left(l\right)}\left(z,a,\eta_{it}^{\left(l\right)}\right)\leq x^{\left(l\right)}\right|Z_{it}=z,A_{i}=a,W_{i}=w\right)$
$=$ $\ \mathbb{P}\left(\left.\eta_{it}^{\left(l\right)}\leq g^{\left(l\right)^{-1}}\left(x^{\left(l\right)},z,a\right)\right|A_{i}=a,W_{i}=w\right)$
$=$ $\ F_{\left.\eta_{it}^{\left(l\right)}\right|A_{i},W_{i}}\left(\left.g^{\left(l\right)^{-1}}\left(x^{\left(l\right)},z,a\right)\right|a,w\right),$ (A.1)
where the first equality holds by $\left.X_{it}\perp A_{i}\right|\left(Z_{it},W_{i}\right)$, the second uses (2.2), the third holds by Assumptions 1 and 3(a), and the last holds by definition of the conditional CDF of $\eta_{it}^{\left(l\right)}$ given $\left(A_{i},W_{i}\right)$. By (2.2), the random variable $\eta_{it}^{\left(l\right)}=g^{\left(l\right)^{-1}}\left(X_{it}^{\left(l\right)},Z_{it},A_{i}\right)$ for each $l$, so that plugging in gives
$V_{it}^{\left(l\right)}\coloneqq F_{\left.X_{it}^{\left(l\right)}\right|Z_{it},W_{i}}\left(\left.X_{it}^{\left(l\right)}\right|Z_{it},W_{i}\right)=F_{\left.\eta_{it}^{\left(l\right)}\right|A_{i},W_{i}}\left(\left.\eta_{it}^{\left(l\right)}\right|A_{i},W_{i}\right),$ (A.2)
which establishes a one-to-one mapping to $\eta_{it}^{\left(l\right)}$ given $A_{i}$ and $W_{i}$ under Assumption 3(b). Let $V_{it}\coloneqq\left(V_{it}^{\left(1\right)},...,V_{it}^{\left(d_{X}\right)}\right)^{{}^{\prime}}$. By (A.2), $V_{it}$ uniquely determines the vector of $\eta_{it}$ given $A_{i}$ and $W_{i}$.
@@
where the first and second equalities hold by the LIE, the third holds because $V_{it}$ is a measurable function of $X_{it},\ Z_{it},$ and $W_{i}$, all of which are also conditioned on, and the fifth equality holds by $\left.\left(X_{it},Z_{it}\right)\perp A_{i}\right|W_{i}$ by Assumption 2.
@@
$\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]=X_{it}^{{}^{\prime}}b_{1}\left(V_{it},W_{i}\right).$ (A.5)
@@
$\mathbb{E}\left[\left.\mathbb{E}\left[\left.X_{it}Y_{it}\right|X_{it},V_{it},W_{i}\right]\right|V_{it},W_{i}\right]=\mathbb{E}\left[\left.X_{it}X_{it}^{{}^{\prime}}\right|V_{it},W_{i}\right]b_{1}\left(V_{it},W_{i}\right),$
@@
$b_{1}\left(V_{it},W_{i}\right)=\left(\mathbb{E}\left[\left.X_{it}X_{it}^{{}^{\prime}}\right|V_{it},W_{i}\right]\right)^{-1}\mathbb{E}\left[\left.X_{it}Y_{it}\right|V_{it},W_{i}\right].$
@@
Given $b_{1}\left(V_{it},W_{i}\right)$, I use the LIE to identify $\overline{b}=\mathbb{E}\beta_{it}$ and $b\left(X_{it}\right)\coloneqq\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ by
$\overline{b}=\mathbb{E}\left[b_{1}\left(V_{it},W_{i}\right)\right]\text{ and }b\left(X_{it}\right)=\mathbb{E}\left[\left.b_{1}\left(V_{it},W_{i}\right)\right|X_{it}\right],$ (A.6)
@@
When ex-post shock $\epsilon_{it}$ is included additively in (2.1) and $\mathbb{E}\epsilon_{it}=0$, equations (A.4) and (A.5) still hold, so the identification result (A.6) holds without any changes. When ex-post shock $\upsilon_{it}$ is also included in $\beta_{it}\coloneqq\beta\left(A_{i},\eta_{it},\upsilon_{it}\right)$, I need to change $b_{2}\left(A_{i},V_{it},W_{i}\right)$ in (A.3) to be $b_{2t}\left(A_{i},V_{it},W_{i}\right)$ because although $\upsilon_{it}$ is independent of all other variables, the PDF of $\upsilon_{it}$ itself may depend on $t$, i.e.,
$\ \mathbb{E}\left[\left.\beta_{it}\right|X_{it},A_{i},V_{it},W_{i}\right]$
$=$ $\ \mathbb{E}\left[\left.\beta\left(A_{i},\varepsilon_{it},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\upsilon_{it}}\right)\right|g\left(Z_{it},A_{i},\eta_{it}\right),A_{i},V_{it},W_{i}\right]$
$=$ $\ \mathbb{E}\left[\left.\beta\left(A_{i},\varepsilon_{it},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\upsilon_{it}}\right)\right|A_{i},V_{it},W_{i}\right]$
$\eqqcolon$ $\ b_{1{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}t}}\left(A_{i},V_{it},W_{i}\right).$ (A.7)
As a result, I need to change $b\left(X_{it}\right)$ to be $b_{t}\left(X_{it}\right)$ in (A.6). The rest of the proof goes through as before.
When exogenous regressors are included in (2.1), I can define $X_{it}=\left(U_{it}^{{}^{\prime}},Z_{1,it}^{{}^{\prime}}\right)^{{}^{\prime}}$ and $Z_{it}=\left(Z_{1,it}^{{}^{\prime}},Z_{2,it}^{{}^{\prime}}\right)^{{}^{\prime}}$, and rewrite model (2.1)–(2.2) to be
$Y_{it}$ $=X_{it}^{{}^{\prime}}\beta\left(A_{i},\varepsilon_{it}\right),$
@@
Then, I replace $X_{it}$ by $U_{it}$ in Assumptions 13 as well as in the definition of $V_{it}$, and the analysis goes through as before. Note that I keep $X_{it}$ in Assumption 4 or 4 because I need to perturb the whole $X_{it}$ vector instead of just endogenous coordinates $U_{it}$ in (A.5) to identify $b_{1}\left(V_{it},W_{i}\right)$. As discussed in footnote 16, the residual variation requirement of Assumption 4 or 4 is easier to be satisfied when exogenous $Z_{1,it}$ is included in $X_{it}$ as the support of $Z_{1,it}$ is unaffected by $V_{it}$ and $W_{i}$. ∎
@@
I denote $\sum_{i=1}^{n}$ by $\sum_{i}$ and omit all $t$-subscripts. As the result for a finite-dimensional vector-valued $\beta$ can be established by proving for each of its coordinates and combining the results using the triangle inequality, I assume $\beta$ is a scalar in this proof. I focus on $\widehat{b}\left(x\right)$, since the result for $\widehat{\overline{b}}$ follows immediately by setting $r^{M_{3}}\left(\cdot\right)\equiv 1$. Let $p_{i}^{m_{2}}\coloneqq p^{m_{2}}\left(V_{i},W_{i}\right)$, $r_{i}\coloneqq r^{M_{3}}\left(X_{i}\right)$, and $\overline{p}_{i}=\overline{p}^{M_{2}}\left(S_{i}\right)$. Following Imbens and Newey (2009), I can normalize $\mathbb{E}p_{i}^{m_{2}}p_{i}^{m_{2}}{}^{{}^{\prime}}=I_{m_{2}}$ by Assumption 5 and $R\coloneqq\mathbb{E}r_{i}r_{i}^{{}^{\prime}}=I_{M_{3}}$ by Assumption 6, which imply $\overline{P}\coloneqq\mathbb{E}\overline{p}_{i}\overline{p}_{i}^{{}^{\prime}}=\mathbb{E}\left[I_{d_{X}}\otimes\left(p_{i}^{m_{2}}p_{i}^{m_{2}}{}^{{}^{\prime}}\right)\right]=I_{M_{2}}$. Furthermore, I have $\lambda_{\text{min}}\left(\widehat{R}\right)\geq C>0$ with probability approaching one by Newey (1997). Let $\widetilde{B}\coloneqq\left(b_{1}\left(\widehat{v}_{1},w_{1}\right),...,b_{1}\left(\widehat{v}_{n},w_{n}\right)\right)^{{}^{\prime}}$.
@@
$\leq$ $\ \left(\widehat{B}-\widetilde{B}\right)^{{}^{\prime}}r\widehat{R}^{-1}r^{{}^{\prime}}\left(\widehat{B}-\widetilde{B}\right)+\left(\widetilde{B}-B\right)^{{}^{\prime}}r\widehat{R}^{-1}r^{{}^{\prime}}\left(\widetilde{B}-B\right)$
$\ +\left(B-B^{X}\right)^{{}^{\prime}}r\widehat{R}^{-1}r^{{}^{\prime}}\left(B-B^{X}\right)+\left(B^{X}-r\rho^{M_{3}}\right)^{{}^{\prime}}r\widehat{R}^{-1}r^{{}^{\prime}}\left(B^{X}-r\rho^{M_{3}}\right),$ (A.8)
where $B\coloneqq\left(b_{1}\left(v_{1},w_{1}\right),...,b_{1}\left(v_{n},w_{n}\right)\right)^{{}^{\prime}}$ and $B^{X}\coloneqq\left(b\left(x_{1}\right),...,b\left(x_{n}\right)\right)^{{}^{\prime}}$. Hence, $\xi=B-B^{X}$. I analyze the RHS of (A.8) term by term.
By Lemma S.5 of Imbens and Newey (2009), $\left\|n^{-1}\sum_{i}\widehat{\overline{p}}_{i}\widehat{\overline{p}}_{i}^{{}^{\prime}}-I\right\|=o_{p}\left(1\right)$. Then,
$\ n^{-2}\left(\widehat{B}-\widetilde{B}\right)^{{}^{\prime}}r\widehat{R}^{-1}r^{{}^{\prime}}\left(\widehat{B}-\widetilde{B}\right)\leq Cn^{-1}\left(\widehat{B}-\widetilde{B}\right)^{{}^{\prime}}\left(\widehat{B}-\widetilde{B}\right)$
$=$ $\ Cn^{-1}\sum_{i}\left(\widehat{\overline{p}}_{i}^{{}^{\prime}}\left(\widehat{\alpha}^{M_{2}}-\alpha^{M_{2}}\right)+\left(\widehat{\overline{p}}_{i}^{{}^{\prime}}\alpha^{M_{2}}-b_{1}\left(\widehat{v}_{i},w_{i}\right)\right)\right)^{2}$
$\leq$ $\ C\left\|\widehat{\alpha}^{M_{2}}-\alpha^{M_{2}}\right\|^{2}+C\sup_{s\in\mathcal{S}}\left\|\overline{p}^{M_{2}}\left(s\right)^{{}^{\prime}}\alpha^{M_{2}}-b_{1}\left(v,w\right)\right\|^{2}=O_{p}\left(\Delta_{2n}^{2}\right),$ (A.9)
where the first inequality holds because $n^{-1}r\widehat{R}^{-1}r^{{}^{\prime}}$ is idempotent, the last inequality holds by the Cauchy-Schwarz inequality (CS) and $\left\|n^{-1}\sum_{i}\widehat{\overline{p}}_{i}\widehat{\overline{p}}_{i}^{{}^{\prime}}-I\right\|=o_{p}\left(1\right)$, and the last equality uses Lemma 1.
@@
$n^{-2}\left(\widetilde{B}-B\right)^{{}^{\prime}}r\widehat{R}^{-1}r^{{}^{\prime}}\left(\widetilde{B}-B\right)\leq$ $\ Cn^{-1}\sum_{i}\left(b_{1}\left(\widehat{v}_{i},w_{i}\right)-b_{1}\left(v_{i},w_{i}\right)\right)^{2}$
@@
$\ n^{-2}\mathbb{E}\left[\left.\left(B-B^{X}\right)^{{}^{\prime}}r\widehat{R}^{-1}r^{{}^{\prime}}\left(B-B^{X}\right)\right|{\bf X}\right]$
$=$ $\ n^{-2}\text{tr}\left\{\mathbb{E}\left[\left.\xi^{{}^{\prime}}r\widehat{R}^{-1}r^{{}^{\prime}}\xi\right|{\bf X}\right]\right\}=n^{-2}\text{tr}\left\{\mathbb{E}\left[\left.\xi\xi^{{}^{\prime}}\right|{\bf X}\right]r\widehat{R}^{-1}r^{{}^{\prime}}\right\}$
$\leq$ $\ n^{-2}\text{tr}\left\{CIr\widehat{R}^{-1}r^{{}^{\prime}}\right\}=Cn^{-1}\text{tr}\left\{\widehat{R}^{-1}\widehat{R}\right\}=Cn^{-1}M_{3},$
@@
$n^{-2}\left(B^{X}-R\rho^{M_{3}}\right)^{{}^{\prime}}r\widehat{R}^{-1}r^{{}^{\prime}}\left(B^{X}-R\rho^{M_{3}}\right)\leq$ $\ n^{-1}\left\|B^{X}-R\rho^{M_{3}}\right\|^{2}=O_{p}\left(M_{3}^{-2d_{3}/j_{3}}\right).$ (A.11)
By $\lambda_{\text{min}}\left(\widehat{R}\right)\geq C>0$ and the result that $\mathbb{E}\left[\left.\left|Y_{n}\right|\right|Z_{n}\right]=O_{p}\left(r_{n}\right)$ implies $Y_{n}=O_{p}\left(r_{n}\right)$ (Conditional Markov’s inequality, henceforth denoted as “CM”),
@@
[…]
$\leq$ $\ 2\left\|\widehat{\rho}^{M_{3}}-\rho^{M_{3}}\right\|^{2}+2\sup_{x\in\mathcal{X}}\left|b\left(x\right)-r^{M_{3}}\left(x\right)^{{}^{\prime}}\rho^{M_{3}}\right|^{2}=O_{p}\left(\Delta_{3n}^{2}\right),$
@@
$\sup_{x\in\mathcal{X}}\left\|\widehat{b}\left(x\right)-b\left(x\right)\right\|$ $\leq\sup_{x\in\mathcal{X}}\left\|r^{M_{3}}\left(x\right)\right\|\left\|\widehat{\rho}^{M_{3}}-\rho^{M_{3}}\right\|+\sup_{x\in\mathcal{X}}\left|b\left(x\right)-r^{M_{3}}\left(x\right)^{{}^{\prime}}\rho^{M_{3}}\right|$
@@
$\Omega_{1}$ $\coloneqq\overline{p}^{M_{2}}\left(v,w\right)^{{}^{\prime}}P^{-1}\left(\Sigma+\Sigma_{1}\right)P^{-1}\overline{p}^{M_{2}}\left(v,w\right),\ \Sigma\coloneqq\mathbb{E}p_{i}p_{i}^{{}^{\prime}}u_{i}^{2},$ (A.13)

### v3

@@ 6 Conclusion
This paper proposes a new TERC model in which regressors are correlated with random coefficients through not only a fixed effect but also a time-varying shock—an empirically relevant feature consistent with optimizing behavior in many applications. I construct feasible control variables for both the fixed effect and the time-varying shock, and use the resulting residual variation in regressors to identify the APE and LAR. I then develop three-step series estimators and establish their convergence rates and asymptotic normality. In an application to Chinese manufacturing, the estimates reveal substantial cross-firm dispersion in output elasticities, part of which is explained by observable firm characteristics.
I propose two directions for future research. First, beyond the low-order moments studied here, policymakers may be interested in the full distribution of random coefficients. One—admittedly demanding—route is to identify moments of all orders by induction and recover the distribution via moment determinacy (Stoyanov (2000)). Second, it remains open whether elements of the approach can be extended to dynamic linear or nonlinear panel models, such as those in Marx, Tamer, and Tang (2024) and Liu, Poirier, and Shiu (2025). Doing so would likely require additional structure governing how lagged outcomes (or state variables) co-move with the random coefficients.
Appendix
Appendix A provides proofs for all theorems and lemmas in the main text, along with additional theoretical results. Appendix B reports additional empirical results. Appendix C presents the simulation results.
Appendix A Proofs and Supplementary Theoretical Results
Appendix A.1 contains proofs of all the lemmas and theorems in the main text. Appendix A.2 introduces Proposition 1, which provides sufficient conditions for Assumption 2 by adapting the nonparametric exchangeability condition of Altonji and Matzkin (2005). Appendix A.3 further discusses Assumptions 2, 4, and 4.
A.1 Proofs
@@
I first prove that $V_{it}$ is a control for $\eta_{it}$ given $A_{i}$ and $W_{i}$. Then, I show $\mathbb{E}\left[\left.\beta_{it}\right|X_{it},V_{it},W_{i}\right]$ does not depend on $X_{it}$ via the law of iterated expectations. Finally, I identify $\mathbb{E}\beta_{it}$ and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ by leveraging the residual variation in $X_{it}$ given $V_{it}$ and $W_{i}$. I present how the inclusion of exogenous shocks $\upsilon_{it}$ and $\widetilde{\epsilon}_{it}$ as well as exogenous coordinates in the regressors affects the analysis at the end of this proof.
Without loss of generality, I assume that $d_{\eta}=d_{X}$ and that each coordinate of $\eta_{it}$ enters the corresponding coordinate of $X_{it}$. [^3] This is without loss of generality because under Assumption 1, I can always redefine $\widetilde{\eta}$ to be the vector that collects the coordinates of $\eta$ that enters each coordinate of $g$ function. By Assumption 2, I have $\left.A_{i}\perp\left(X_{it},Z_{it}\right)\right|W_{i}$, which implies $\left.X_{it}\perp A_{i}\right|\left(Z_{it},W_{i}\right)$. Thus, for each $l\in\left\{1,\ldots,d_{X}\right\}$ and any on-support $\left(x_{l},z,a,w\right)$, I have
$\ F_{\left.X_{it,l}\right|Z_{it},W_{i}}\left(\left.x_{l}\right|z,w\right)$
$=$ $\ F_{\left.X_{it,l}\right|Z_{it},A_{i},W_{i}}\left(\left.x_{l}\right|z,a,w\right)$
$=$ $\ \mathbb{P}\left(\left.g_{l}\left(z,a,\eta_{it,l}\right)\leq x_{l}\right|Z_{it}=z,A_{i}=a,W_{i}=w\right)$
$=$ $\ \mathbb{P}\left(\left.\eta_{it,l}\leq g_{l}^{-1}\left(x_{l},z,a\right)\right|A_{i}=a,W_{i}=w\right)$
$=$ $\ F_{\left.\eta_{it,l}\right|A_{i},W_{i}}\left(\left.g_{l}^{-1}\left(x_{l},z,a\right)\right|a,w\right),$ (A.1)
where the first equality holds by $\left.X_{it}\perp A_{i}\right|\left(Z_{it},W_{i}\right)$, the second uses (2.2), the third holds by Assumptions 1 and 3(a), and the last holds by definition. By (2.2), the random variable $\eta_{it,l}=g_{l}^{-1}\left(X_{it,l},Z_{it},A_{i}\right)$ for each $l$, so that plugging in gives
$V_{it,l}\coloneqq F_{\left.X_{it,l}\right|Z_{it},W_{i}}\left(\left.X_{it,l}\right|Z_{it},W_{i}\right)=F_{\left.\eta_{it,l}\right|A_{i},W_{i}}\left(\left.\eta_{it,l}\right|A_{i},W_{i}\right),$ (A.2)
which establishes a one-to-one mapping to $\eta_{it,l}$ given $A_{i}$ and $W_{i}$ under Assumption 3(b). Let $V_{it}\coloneqq(V_{it,1},\dots,V_{it,d_{X}})^{\prime}$. By (A.2), $V_{it}$ uniquely determines the vector of $\eta_{it}$ given $A_{i}$ and $W_{i}$.
@@
where the first and second equalities hold by the law of iterated expectations, the third holds because $V_{it}$ is a measurable function of $X_{it},\ Z_{it},$ and $W_{i}$, all of which are also conditioned on, and the fifth equality holds by $\left.\left(X_{it},Z_{it}\right)\perp A_{i}\right|W_{i}$ by Assumption 2.
@@
$\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]=X_{it}^{\prime}b_{1}\left(V_{it},W_{i}\right).$ (A.5)
@@
$\mathbb{E}\left[\left.\mathbb{E}\left[\left.X_{it}Y_{it}\right|X_{it},V_{it},W_{i}\right]\right|V_{it},W_{i}\right]=\mathbb{E}\left[\left.X_{it}X_{it}^{\prime}\right|V_{it},W_{i}\right]b_{1}\left(V_{it},W_{i}\right),$
@@
$b_{1}\left(V_{it},W_{i}\right)=\left(\mathbb{E}\left[\left.X_{it}X_{it}^{\prime}\right|V_{it},W_{i}\right]\right)^{-1}\mathbb{E}\left[\left.X_{it}Y_{it}\right|V_{it},W_{i}\right].$
@@
Given $b_{1}\left(V_{it},W_{i}\right)$, I use the law of iterated expectations to identify $\overline{b}$ and $b(X_{it})$ by
$\overline{b}=T^{-1}\sum_{t=1}^{T}\mathbb{E}\left[b_{1}\left(V_{it},W_{i}\right)\right]\text{ and }b\left(X_{it}\right)=\mathbb{E}\left[\left.b_{1}\left(V_{it},W_{i}\right)\right|X_{it}\right],$ (A.6)
@@
When ex-post shock $\widetilde{\epsilon}_{it}$ is included additively in (2.1) and $\mathbb{E}\widetilde{\epsilon}_{it}=0$, equations (A.4) and (A.5) still hold, so the identification result (A.6) holds without any changes. When ex-post shock $\upsilon_{it}$ is also included in $\beta_{it}\coloneqq\beta\left(A_{i},\varepsilon_{it},\upsilon_{it}\right)$, since $b(X_{it})$ is already a time-varying function of $X_{it}$ due to the distribution of $(\varepsilon_{it},\eta_{it})$ being possibly time-varying, adding $\upsilon_{it}$ which is independent of everything else does not affect the identification proof.
Finally, when exogenous regressors are included in (2.1), I let $X_{it}=\left(U_{it}^{\prime},Z_{it,1}^{\prime}\right)^{\prime}$ and $Z_{it}=\left(Z_{it,1}^{\prime},Z_{it,2}^{\prime}\right)^{\prime}$, and rewrite model (2.1)–(2.2) to be
$Y_{it}$ $=X_{it}^{\prime}\beta\left(A_{i},\varepsilon_{it}\right),$
@@
Then, I replace $X_{it}$ by $U_{it}$ in Assumptions 13 as well as in the definition of $V_{it}$, and the analysis goes through as before. Note that I keep $X_{it}$ in Assumption 4 or 4 because I need to perturb the whole $X_{it}$ vector instead of just endogenous coordinates $U_{it}$ in (A.5) to identify $b_{1}\left(V_{it},W_{i}\right)$. As discussed in Appendix A.3, the residual variation requirement of Assumption 4 or 4 is easier to be satisfied when exogenous $Z_{it,1}$ is included in $X_{it}$ as the support of $Z_{it,2}$ is unaffected by $V_{it}$ and $W_{i}$. ∎
@@
I denote $\sum_{i=1}^{n}$ by $\sum_{i}$ and omit all $t$-subscripts. As the result for a finite-dimensional vector-valued $\beta$ can be established by proving for each of its coordinates and combining the results using the triangle inequality, I assume $\beta$ is a scalar in this proof. I focus on $\widehat{b}\left(x\right)$, since the result for $\widehat{\overline{b}}$ follows immediately by setting $r^{M_{3}}\equiv 1$. Let $p_{i}^{m_{2}}\coloneqq p^{m_{2}}\left(V_{i},W_{i}\right)$, $r_{i}\coloneqq r^{M_{3}}\left(X_{i}\right)$, and $\overline{p}_{i}=\overline{p}^{M_{2}}\left(S_{i}\right)$. Following Imbens and Newey (2009), I can normalize $\mathbb{E}p_{i}^{m_{2}}p_{i}^{m_{2}}{}^{\prime}=I_{m_{2}}$ by Assumption 5 and $R\coloneqq\mathbb{E}r_{i}r_{i}^{\prime}=I_{M_{3}}$ by Assumption 6, which imply $\overline{P}\coloneqq\mathbb{E}\overline{p}_{i}\overline{p}_{i}^{\prime}=\mathbb{E}\left[I_{d_{X}}\otimes\left(p_{i}^{m_{2}}p_{i}^{m_{2}}{}^{\prime}\right)\right]=I_{M_{2}}$. Furthermore, I have $\lambda_{\text{min}}\left(\widehat{R}\right)\geq C>0$ with probability approaching one by Newey (1997). Let $\widetilde{B}\coloneqq\left(b_{1}\left(\widehat{v}_{1},w_{1}\right),\ldots,b_{1}\left(\widehat{v}_{n},w_{n}\right)\right)^{\prime}$.
@@
$\leq$ $\ \left(\widehat{B}-\widetilde{B}\right)^{\prime}r\widehat{R}^{-1}r^{\prime}\left(\widehat{B}-\widetilde{B}\right)+\left(\widetilde{B}-B\right)^{\prime}r\widehat{R}^{-1}r^{\prime}\left(\widetilde{B}-B\right)$
$\ +\left(B-B^{X}\right)^{\prime}r\widehat{R}^{-1}r^{\prime}\left(B-B^{X}\right)+\left(B^{X}-R\rho^{M_{3}}\right)^{\prime}r\widehat{R}^{-1}r^{\prime}\left(B^{X}-r\rho^{M_{3}}\right),$ (A.7)
where $B\coloneqq\left(b_{1}\left(v_{1},w_{1}\right),\ldots,b_{1}\left(v_{n},w_{n}\right)\right)^{\prime}$ and $B^{X}\coloneqq\left(b\left(x_{1}\right),\ldots,b\left(x_{n}\right)\right)^{\prime}$. Hence, $\xi=B-B^{X}$. I analyze the RHS of (A.7) term by term.
By Lemma S.5 of Imbens and Newey (2009), $\left\|n^{-1}\sum_{i}\widehat{\overline{p}}_{i}\widehat{\overline{p}}_{i}^{\prime}-I\right\|=o_{p}\left(1\right)$. Then,
[…]
$=$ $\ Cn^{-1}\sum_{i}\left(\widehat{\overline{p}}_{i}^{\prime}\left(\widehat{\alpha}^{M_{2}}-\alpha^{M_{2}}\right)+\left(\widehat{\overline{p}}_{i}^{\prime}\alpha^{M_{2}}-b_{1}\left(\widehat{v}_{i},w_{i}\right)\right)\right)^{2}$
$\leq$ $\ C\left\|\widehat{\alpha}^{M_{2}}-\alpha^{M_{2}}\right\|^{2}+C\sup_{s\in\mathcal{S}}\left\|\overline{p}^{M_{2}}\left(s\right)^{\prime}\alpha^{M_{2}}-b_{1}\left(v,w\right)\right\|^{2}=O_{p}\left(\Delta_{2n}^{2}\right),$ (A.8)
where the first inequality holds because $n^{-1}r\widehat{R}^{-1}r^{\prime}$ is idempotent, the last inequality holds by the Cauchy–Schwarz inequality (CS) and $\left\|n^{-1}\sum_{i}\widehat{\overline{p}}_{i}\widehat{\overline{p}}_{i}^{\prime}-I\right\|=o_{p}\left(1\right)$, and the last equality uses Lemma 1.
@@
$n^{-2}\left(\widetilde{B}-B\right)^{\prime}r\widehat{R}^{-1}r^{\prime}\left(\widetilde{B}-B\right)\leq$ $\ Cn^{-1}\sum_{i}\left(b_{1}\left(\widehat{v}_{i},w_{i}\right)-b_{1}\left(v_{i},w_{i}\right)\right)^{2}$
@@
$\ n^{-2}\mathbb{E}\left[\left.\left(B-B^{X}\right)^{\prime}r\widehat{R}^{-1}r^{\prime}\left(B-B^{X}\right)\right|{\bf X}\right]$
$=$ $\ n^{-2}\text{tr}\left\{\mathbb{E}\left[\left.\xi^{\prime}r\widehat{R}^{-1}r^{\prime}\xi\right|{\bf X}\right]\right\}=n^{-2}\text{tr}\left\{\mathbb{E}\left[\left.\xi\xi^{\prime}\right|{\bf X}\right]r\widehat{R}^{-1}r^{\prime}\right\}$
$\leq$ $\ n^{-2}\text{tr}\left\{CIr\widehat{R}^{-1}r^{\prime}\right\}=Cn^{-1}\text{tr}\left\{\widehat{R}^{-1}\widehat{R}\right\}=Cn^{-1}M_{3},$
@@
$n^{-2}\left(B^{X}-R\rho^{M_{3}}\right)^{\prime}r\widehat{R}^{-1}r^{\prime}\left(B^{X}-R\rho^{M_{3}}\right)\leq$ $\ n^{-1}\left\|B^{X}-R\rho^{M_{3}}\right\|^{2}=O_{p}\left(M_{3}^{-2d_{3}/j_{3}}\right).$ (A.10)
By $\lambda_{\text{min}}\left(\widehat{R}\right)\geq C>0$ and conditional Markov’s (CM) inequality (i.e., $\mathbb{E}\left[\left.\left|Y_{n}\right|\right|Z_{n}\right]=O_{p}\left(r_{n}\right)$ implies $Y_{n}=O_{p}\left(r_{n}\right)$),
@@
$\int\left\|\widehat{b}\left(x\right)-b\left(x\right)\right\|^{2}dF\left(x\right)\leq$ $\int\left(r^{M_{3}}\left(x\right)^{\prime}\left(\widehat{\rho}^{M_{3}}-\rho^{M_{3}}\right)+\left(r^{M_{3}}\left(x\right)^{\prime}\rho^{M_{3}}-b\left(x\right)\right)\right)^{2}dF\left(x\right)$
$\leq$ $\ 2\left\|\widehat{\rho}^{M_{3}}-\rho^{M_{3}}\right\|^{2}+2\sup_{x\in\mathcal{X}}\left|b\left(x\right)-r^{M_{3}}\left(x\right)^{\prime}\rho^{M_{3}}\right|^{2}=O_{p}\left(\Delta_{3n}^{2}\right),$
@@
$\sup_{x\in\mathcal{X}}\left\|\widehat{b}\left(x\right)-b\left(x\right)\right\|$ $\leq\sup_{x\in\mathcal{X}}\left\|r^{M_{3}}\left(x\right)\right\|\left\|\widehat{\rho}^{M_{3}}-\rho^{M_{3}}\right\|+\sup_{x\in\mathcal{X}}\left|b\left(x\right)-r^{M_{3}}\left(x\right)^{\prime}\rho^{M_{3}}\right|$
@@
$\Omega_{1}$ $\coloneqq\overline{p}^{M_{2}}\left(v,w\right)^{\prime}P^{-1}\left(\Sigma+\Sigma_{1}\right)P^{-1}\overline{p}^{M_{2}}\left(v,w\right),\ \Sigma\coloneqq\mathbb{E}p_{i}p_{i}^{\prime}u_{i}^{2},$ (A.12)